# Unlocking State-Tracking in Linear RNNs Through Negative Eigenvalues

### v1

@@ C.3 Krohn-Rhodes Theorem
Before presenting the proof for Theorem 4, we provide the statement for the landmark result of Krohn-Rhodes (Krohn & Rhodes, 1965), after giving the definition of cascade product of two FSA.
@@ C.4 Proof of Theorem 4
We apply the Krohn-Rhodes theorem (Theorem 5) to write $\mathcal{A}$ as the cascade product FSA $\mathcal{C}=\mathcal{A}^{(s)}\circ\cdots\circ\mathcal{A}^{(1)}$ with each FSA $\mathcal{A}^{(i)}=\big{(}\Sigma^{(i)},Q^{(i)},q_{0}^{(i)},\delta^{(i)}\big{)}$ being permutation-reset and we show how the LRNN can implement $\mathcal{C}$ by first showing how its $i$-th layer, with the structure in 1, can implement $\mathcal{A}^{(i)}$.
Let $n=|Q^{(i)}|$ and without loss of generality assume that $\Sigma=\{1,2,\dots,|\Sigma|\}$ and $Q^{(i)}=\{1,2,\dots,n\}$ with $q_{0}^{(i)}=1$. For every $w\in\Sigma^{(i)}$ we set ${\bm{A}}^{(i)}(w)\in\{0,1\}^{n\times n}$, ${\bm{B}}^{(i)}(w)\in\{0,1\}^{n}$ such that $q,q^{\prime}\in Q^{(i)}$
@@
where we removed the nested parenthesis for simplicity. Hence, the first $s$ elements of $y_{t}^{(s)}$ are exactly the output of the cascade FSA $\mathcal{C}$. Note that our construction can be implemented in finite precision since we only used matrices/vectors with entries either in $\{0,1\}$, requiring only one bit, or in $Q^{(i)}\subset\mathbb{N}$, that can also be implemented using finite precision with $|Q^{(i)}|$ integers, requiring $\log_{2}(|Q^{(i)}|)$ bits. Note that we can exclude the last element of $y_{t}^{(s)}$ by changing $\mathrm{dec}^{(s)}$, to get a width of $\mathbb{N}^{s}$.
@@
It follows that $\mathcal{A}^{(i)}(w)\in\mathcal{M}_{n}^{n}([-1,1])$ for $i\in\{1,\dots,s\}$. ∎
@@ Appendix D LRNNs Can Do Modular Addition Using Only Reflections
In this section, we explain how an LRNN with two layers and using only Householder state transition matrices (reflections) can compute addition modulo $m\in\mathbb{N}$, i.e it can map words $x_{1},\dots,x_{t}$ with $x_{i}\in\{0,\dots,m-1\}$ into $y_{t}=(\sum_{i=1}^{m}x_{i})\bmod m$. This corresponds to solving the group word problem associated with the cyclic group $\mathbb{Z}_{m}$. We note that our modification of DeltaNet, namely DeltaNet [-1,1] can therefore solve addition modulo $m$ with 2 layers.
If the state transition matrices can be a generic rotation matrices, then a LRNN can perform addition modulo $m$ using just one layer by mapping each element of $\mathbb{Z}_{m}$ to the corresponding $2\times 2$ rotation matrix as shown in Section C.2. Such construction requires a number of states for the LRNN equal to $m$, i.e. the number of elements of the group $\mathbb{Z}_{m}$. However, since we assume that state transition matrices are reflections, we cannot map each element of the group to a rotation (since those are a product of 2 reflections) and our construction for the LRNN will require two layers. Specifically, the first layer will count modulo $2$, i.e. it will output the sequence ${\bm{y}}^{(1)}_{1},\dots,{\bm{y}}^{(1)}_{t}$ where ${\bm{y}}^{(1)}_{i}=(x_{i},i\bmod 2)$, while the second layer will have $2m$ states and will use two different reflection matrices for each group element, depending on the value of $y^{(1)}_{i,2}=i\bmod 2$. Formally, we have the following result.
@@
An LRNN with two layers in the form (1), where ${\bm{A}}:\mathbb{N}\to\{-1\}$ for the first layer and ${\bm{A}}:\mathbb{R}^{2}\to\mathcal{M}_{1}^{2}(\{-1\})$ for the second layer, with $\mathcal{M}_{1}^{2}$ is defined in (5), can perform addition modulo $m$. In particular, the LRNN will have 2 scalar states in the first layer and $2m$ states, each being a vector in $\mathbb{R}^{2}$, in the second layer.
@@
$\theta(i,1)=\frac{(1-2i)\pi}{m},\quad\theta(i,0)=\frac{(2i+1)\pi}{m},\quad\text{ for all }i\in\{0,\dots,m-1\}.$
@@ E.1.1 Details on the experimental setup
Like Beck et al. (2024), we trained each model with sequence lengths ranging from 3 to 40 and evaluated on lengths from 40 to 256, to understand the length generalization capabilities. We compared mLSTM and sLSTM with two models: Mamba (Gu & Dao, 2023) and DeltaNet (Yang et al., 2024b). Moreover, we also include a Transformer (Vaswani et al., 2017) baseline. For parity, all models contain 2 blocks layers), with 4 heads for the xLSTM and DeltaNet models. We set the embedding and heads’ dimensions to 128. For Mamba and DeltaNet, we also enable the 1-D depthwise-separable convolution layer with kernel size equal to 4 after the query/key/value projection. For modular arithmetic, we increase the number of layers to 3 and use a gradient clipping norm of 1.0 for Transformer, Mamba, and DeltaNet, while for mLSTM and sLSTM we decrease the embedding size and number of heads to 64 and 1, respectively, as well as use a standard initialization for the bias parameters. We train each model using AdamW (Loshchilov & Hutter, 2019) without gradient clipping, using 3 different learning rates (1e-2, 1e-3, 5e-4 1e-4), with 3 different seeds each. We pick the best based on the median of the 3 seeds for every learning rate value. We use a batch size of 1024 (except for mLSTM, where we use 512 due to OOM error) and a cosine annealing learning rate schedule (Loshchilov & Hutter, 2017) (minimum learning rate: 1e-6) after 10% warm-up steps. The weight decay is set to 0.1 during training. We train on every task for 100k steps in total. At each training step, we make sure to generate a valid random sample from the task at hand (see below).
@@ E.1.2 Details on the evaluated tasks
In Section 5.1 we conducted empirical evaluations on 3 tasks –namely parity, modular arithmetic without brackets and with brackets – from various levels of the Chomsky Hierarchy, as proposed by Deletang et al. (2023) and similarly used in xLSTM (Beck et al., 2024). Details for each task are given below, where $|\Sigma|$ is the vocabulary size and $Acc_{rand}$ is the accuracy of random guessing:
@@
  • Modular Arithmetic w/o Brackets ($|\Sigma|=10$, $Acc_{rand}=1/(|\Sigma|-5)$). Given a set of special tokens $\Sigma_{s}=\{\mathtt{+,-,*,=,[PAD]}\}$ and a modulus $m\geq 1$, we compute the remainder $y_{t}={\bm{x}}\ mod\ m$, where ${\bm{x}}=x_{1}\dots x_{t}\in\Sigma^{t}$ and $y_{t}\in\{1,\dots,m-1\}$. Here, $\Sigma=\Sigma_{s}\cup\{0,\dots,m-1\}$. In our experiments $m=5$. An example sequence is as follows:
@@
  • Modular Arithmetic w/ Brackets ($|\Sigma|=12$, $Acc_{rand}=1/(|\Sigma|-7)$). Same definition as the modular arithmetic without brackets with a set of special tokens $\Sigma_{s}=\{\mathtt{+,-,*,=,),(,[PAD]}\}$. In our experiments $m=5$. An example sequence is as follows:
@@ E.2.1 Details of the Experiments
  • $\mathbf{S_{5}}$ We sample $x_{i}$ uniformly at random from $\{0,\dots,119\}$. $y_{i}$ is computed as the product of the permutations corresponding to $x_{1},\dots,x_{i}$.
@@
For each input, we also add a beginning of sequence token. For each setup, we always sample $1.6$M examples for training and $40K$ examples of length $500$ for testing. We note that we are using a substantially larger training set compared to (Merrill & Sabharwal, 2023), to reduce the chances of overfitting. We run 3 seeds for each method, changing the network initialization and sampling of the minibatches. The train and validation datasets are kept the same across runs.
We train all models using AdamW with weight decay $0.01$, learning rate $0.0001$, gradient clipping to $1.0$, and a batch size of $512$.
Both DeltaNet and Mamba models use an embedding dimension of $128$ and $4$ heads for DeltaNet. In the case of DeltaNet, we do not use the $1$-D convolutions for these experiments. Other parameters are kept as default.
@@
where $n$ is set to $32$ for efficiency reasons (memory and compute time grow quickly with $n$). After that, we flatten each ${\bm{H}}_{i}$ into a vector and apply first a projection on the unit ball and then a linear decoder to get the final outputs. The projection was added to increase stability since we do not bound the norm of ${\bm{A}}(x_{i})$. Since this model uses a full matrix, with $n\geq 5$ it should be fully able to learn $S_{5}$ without restricting the transitions in input or using more tokens per transition. However, in some situations, the performance degrades quickly after some input sequence length, probably because the norm of the learned ${\bm{A}}(x_{i})$ is not close enough to one.
Plots with all runs. We report the plots with all 3 runs per method in Figure 5 (In Figure 2 we reported only the best one for each method). Despite our efforts to reduce randomness in the training by increasing training time and dataset size, we report that there is still some variability. For example, one of the runs of DeltaNet $[-1,1]$ (5L) on $S_{5}$ with 4 tokens per transition did not manage to learn the task fully.
@@ E.3.2 Details on the evaluated tasks
(a) DeltaNet
(b) Mamba
Figure 7: Learning curves of Mamba (370M) and DeltaNet (340M) when training on 32B tokens of Fine-Web 100B.
@@ E.4 Implementation
We build on the original code for Mamba [^4] https://github.com/state-spaces/mamba and DeltaNet [^5] https://github.com/sustcsonglin/flash-linear-attention. For DeltaNet, implementing the extended eigenvalue range is straightforward, since there is no need to modify the Triton kernel. However, Mamba requires modifications to the CUDA code of the associative scan for both forward and backward passes which however had no impact on computational cost. We ensured the accuracy of the modifications by comparing the results with a naive implementation using a for-loop. For initial testing of the extended eigenvalue range, we used the pure PyTorch implementation of Mamba by Torres (2024). We provide listings of the necessary code changes in Mamba and DeltaNet in Section E.4.1. For DeltaNet, this changes also ${\bm{B}}(x_{t})$ in Table 1 by multiplying it by $2$.
Products in Log-space We note that some diagonal models such as Mamba2 (Dao & Gu, 2024), GLA (Yang et al., 2024a), mLSTM (Beck et al., 2024) take advantage of the fact that all values of the state-transition matrices are positive to compute their repeated products in log-space. Our change would not allow us to do this directly, and early tests on the chunkwise parallel form of GLA showed degraded performance. Therefore, for this work, we decided to focus on Mamba and DeltaNet since they do not compute the products in log space. We mention however, that at the cost of increased computation time, it would be possible to do products in log space by converting each value in the diagonal state-transition matrix to the product of its absolute value and sign. This way, absolute values can be multiplied in log space, while products of signs are coincidentally equivalent to addition modulo 2, i.e. parity, and hence can be done stably. We leave the investigation of this approach to future work. Furthermore, we also believe that our change may be less suited for methods that use a normalized RNN state, such as mLSTM.

### v2

@@ C.3 Krohn-Rhodes Theorem
Before presenting the proof for Theorem 4, we provide the statement for the landmark result of Krohn-Rhodes (Krohn & Rhodes, 1965), after giving the definition of the cascade product of two FSA.
@@ C.4 Proof of Theorem 4
We apply the Krohn-Rhodes theorem (Theorem 5) to write $\mathcal{A}$ as the cascade product FSA $\mathcal{C}=\mathcal{A}^{(s)}\circ\cdots\circ\mathcal{A}^{(1)}$ with each FSA $\mathcal{A}^{(i)}=\big{(}\Sigma^{(i)},Q^{(i)},q_{0}^{(i)},\delta^{(i)}\big{)}$ being permutation-reset and we show how the LRNN can implement $\mathcal{C}$ by first showing how its $i$-th layer, with the structure in (1), can implement $\mathcal{A}^{(i)}$.
Let $n=|Q^{(i)}|$ and without loss of generality assume that $\Sigma=\{1,2,\dots,|\Sigma|\}$ and $Q^{(i)}=\{1,2,\dots,n\}$ with $q_{0}^{(i)}=1$. For every $w\in\Sigma^{(i)}$ we set ${\bm{A}}^{(i)}(w)\in\{0,1\}^{n\times n}$, ${\bm{B}}^{(i)}(w)\in\{0,1\}^{n}$ such that for every $q,q^{\prime}\in Q^{(i)}$
@@
where we removed the nested parenthesis for simplicity. Hence, the first $s$ elements of $y_{t}^{(s)}$ are exactly the output of the cascade FSA $\mathcal{C}$. Note that our construction can be implemented in finite precision since we only used matrices/vectors with entries either in $\{0,1\}$, requiring only one bit, or in $Q^{(i)}\subset\mathbb{N}$, that can also be implemented using finite precision with $|Q^{(i)}|$ integers, requiring $\log_{2}(|Q^{(i)}|)$ bits. Note that we can exclude $w_{t}$ from the output $y_{t}^{(s)}$ by changing $\mathrm{dec}^{(s)}$, to bring the dimension of the output, end hence the width of the LRNN, to $\mathbb{N}^{s}$.
@@
It follows that $\mathcal{A}^{(i)}(w)\in\mathcal{M}_{n}^{n}([-1,1])$ for every $i\in\{1,\dots,s\}$ and $w\in\Sigma^{(i)}$. ∎
@@ Appendix D LRNNs Can Do Modular Addition Using Only Reflections
In this section, we explain how an LRNN with two layers and using only Householder state transition matrices (reflections) can compute addition modulo $m\in\mathbb{N}$, i.e it can map words $x_{1},\dots,x_{t}$ with $x_{i}\in\{0,\dots,m-1\}$ into $y_{t}=(\sum_{i=1}^{m}x_{i})\bmod m$ for arbitrary $t\in\mathbb{N}$. This corresponds to solving the group word problem associated with the cyclic group $\mathbb{Z}_{m}$. We note that our modification of DeltaNet, namely DeltaNet [-1,1] can therefore solve addition modulo $m$ with 2 layers.
If the state transition matrices can be generic rotation matrices, then a LRNN can perform addition modulo $m$ using just one layer by mapping each element of $\mathbb{Z}_{m}$ to the corresponding $2\times 2$ rotation matrix as shown in Section C.2. Such construction requires a number of states for the LRNN equal to $m$, i.e. the number of elements of the group $\mathbb{Z}_{m}$. However, since here we assume that state transition matrices are reflections, we cannot map each element of the group to a rotation (since those are a product of 2 reflections) and our construction for the LRNN will require two layers. Specifically, the first layer will count modulo $2$, i.e. it will output the sequence ${\bm{y}}^{(1)}_{1},\dots,{\bm{y}}^{(1)}_{t}$ where ${\bm{y}}^{(1)}_{i}=(x_{i},i\bmod 2)$, while the second layer will have $2m$ states and will use two different reflection matrices for each group element, depending on the value of $y^{(1)}_{i,2}=i\bmod 2$. Formally, we have the following result.
@@
An LRNN with two layers in the form (1), where ${\bm{A}}:\mathbb{N}\to\{-1\}$ for the first layer and ${\bm{A}}:\mathbb{R}^{2}\to\mathcal{M}_{1}^{2}(\{-1\})$ for the second layer, with $\mathcal{M}_{1}^{2}$ defined in (5), can perform addition modulo $m$. In particular, the LRNN will have 2 scalar states in the first layer and $2m$ states, each being a vector in $\mathbb{R}^{2}$, in the second layer.
@@
$\theta(i,1)=\frac{(1-2i)\pi}{m},\quad\theta(i,0)=\frac{(1+2i)\pi}{m},\quad\text{ for all }i\in\{0,\dots,m-1\}.$
@@ E.1.1 Details on the experimental setup
Like Beck et al. (2024), we trained each model with sequence lengths ranging from 3 to 40 and evaluated on lengths from 40 to 256, to understand the length generalization capabilities. We compared mLSTM and sLSTM with two models: Mamba (Gu & Dao, 2023) and DeltaNet (Yang et al., 2024b). Moreover, we also include a Transformer (Vaswani et al., 2017) baseline. For parity, all models contain 2 blocks (layers), with 4 heads for the xLSTM and DeltaNet models. We set the embedding and heads’ dimensions to 128. For Mamba and DeltaNet, we also enable the 1-D depthwise-separable convolution layer with kernel size equal to 4 after the query/key/value projection. For modular arithmetic, we increase the number of layers to 3 and use a gradient clipping norm of 1.0 for Transformer, Mamba, and DeltaNet, while for mLSTM and sLSTM we decrease the embedding size and number of heads to 64 and 1, respectively, as well as use a standard initialization for the bias parameters. We train each model using AdamW (Loshchilov & Hutter, 2019) without gradient clipping, using 3 different learning rates (1e-2, 1e-3, 5e-4 1e-4), with 3 different seeds each. We pick the best based on the median of the 3 seeds for every learning rate value. We use a batch size of 1024 (except for mLSTM, where we use 512 due to OOM error) and a cosine annealing learning rate schedule (Loshchilov & Hutter, 2017) (minimum learning rate: 1e-6) after 10% warm-up steps. The weight decay is set to 0.1 during training. We train on every task for 100k steps in total. At each training step, we make sure to generate a valid random sample from the task at hand (see below).
@@ E.1.2 Details on the evaluated tasks
In Section 5.1 we conducted empirical evaluations on 3 tasks –parity, modular arithmetic without brackets and with brackets – from various levels of the Chomsky Hierarchy, as proposed by Deletang et al. (2023) and similarly used in xLSTM (Beck et al., 2024). Details for each task are given below, where $|\Sigma|$ is the vocabulary size and $Acc_{rand}$ is the accuracy of random guessing:
@@
  • Modular Arithmetic w/o Brackets ($|\Sigma|=10$, $Acc_{rand}=1/5$). Given a set of special tokens $\Sigma_{s}=\{\mathtt{+,-,*,=,[PAD]}\}$ and a modulus $m\geq 1$, we set $\Sigma=\Sigma_{s}\cup\{0,\dots,m-1\}$ and $y_{t}$ is equal to the result of the operations modulo $m$ in the sequence ${\bm{x}}={\bm{x}}_{1},\dots,{\bm{x}}_{t}$ with $x_{i}\in\Sigma$. In our experiments $m=5$. An example sequence is as follows:
@@
  • Modular Arithmetic w/ Brackets ($|\Sigma|=12$, $Acc_{rand}=1/5$). Same definition as the modular arithmetic without brackets with a set of special tokens $\Sigma_{s}=\{\mathtt{+,-,*,=,),(,[PAD]}\}$. In our experiments $m=5$. An example sequence is as follows:
@@ E.2.1 Details of the Experiments
  • $\mathbf{S_{5}}$ We sample $x_{i}$ uniformly at random from $\{0,\dots,119\}$. $y_{i}$ is computed as the product of the permutations corresponding to $x_{1},\dots,x_{i}$ applied in order from $1$ to $i$.
@@
For each setup, we randomly sample $1.6$M examples for and $40K$ examples of length $500$ to construct the train and test dataset. We note that we are using a substantially larger training set compared to (Merrill & Sabharwal, 2023), to reduce the chances of overfitting. We run 3 seeds for each method, changing the network initialization and sampling of the minibatches. The train and validation datasets are kept the same across runs.
We train all models using AdamW with weight decay $0.01$, learning rate $0.0001$, gradient clipping to $1.0$, and a batch size of $512$. Both DeltaNet and Mamba models use an embedding dimension of $128$ and $4$ heads for DeltaNet. In the case of DeltaNet, we do not use $1$-D convolutions for these experiments. Other parameters are kept as default.
@@
where $n$ is set to $32$ for efficiency reasons (memory and compute time grow quickly with $n$). After that, we flatten each ${\bm{H}}_{i}$ into a vector and apply first a projection on the unit ball and then a linear decoder to get the final outputs. The projection was added to increase stability since we do not bound the norm of ${\bm{A}}(x_{i})$. Since this model uses a full matrix, with $n\geq 5$ it should be fully able to learn $S_{5}$ without restricting the transitions in input or using more tokens per transition. However, in some situations, the performance degrades quickly after some input sequence length, probably because the norm of the learned ${\bm{A}}(x_{i})$ is not close enough to one and hence part of the state either vanish or explode for long sequences.
Plots with all runs. We report the plots with all 3 runs per method in Figure 5 (In Figure 2 we reported only the best one for each method). Despite our efforts to decrease the variance of the results by increasing training time and dataset size, we report that there is still some variability. For example, one of the runs of DeltaNet $[-1,1]$ (5L) on $S_{5}$ with 4 tokens per transition did not achieve a good accuracy.
@@ E.3.1 Details on the experimental setup
Figure 7: Learning curves of DeltaNet 340M (top left), Mamba 370M (top right) and DeltaNet 1.3B (bottom), training on 100B tokens of Fine-Web 100B. 1.3B runs required only 50k optimizer steps versus the 200k of the 340M runs due to the 4x larger batch size. All models trained stably with the same hyperparameters. Training curves were smoothed with a rolling window of 500 steps.
@@ E.4 Implementation
We build on the original code for Mamba [^5] https://github.com/state-spaces/mamba and DeltaNet [^6] https://github.com/sustcsonglin/flash-linear-attention. For DeltaNet, implementing the extended eigenvalue range is straightforward, since there is no need to modify the Triton kernel. However, Mamba requires modifications to the CUDA code of the associative scan for both forward and backward passes which however had no impact on computational cost. We ensured the accuracy of the modifications by comparing the results with a naive implementation using a for-loop. For initial testing of the extended eigenvalue range, we used the pure PyTorch implementation of Mamba by Torres (2024). We provide listings of the necessary code changes in Mamba and DeltaNet in Section E.4.1. For DeltaNet, this changes also ${\bm{B}}(x_{t})$ in Table 1, multiplying it by $2$.
Products in Log-space We note that some diagonal models such as Mamba2 (Dao & Gu, 2024), GLA (Yang et al., 2024a), and mLSTM (Beck et al., 2024) take advantage of the fact that all values of the state-transition matrices are positive to compute their repeated products in log-space. Our change would not allow us to do this directly, and early tests on the chunkwise parallel form of GLA showed degraded performance. Therefore, for this work, we decided to focus on Mamba and DeltaNet since they do not compute the products in log space. We mention however, that at the cost of increased computation time, it would be possible to do products in log space by converting each value in the diagonal state-transition matrix to the product of its absolute value and sign. This way, absolute values can be multiplied in log space, while products of signs are coincidentally equivalent to addition modulo 2, i.e. parity, and hence can be done stably. We leave the investigation of this approach to future work. Furthermore, we also believe that our change may be less suited for methods that use a normalized RNN state, such as mLSTM, since it might happen that the normalization term can be very close to zero due to the negative values.